Bent Gregorian Technical Observing with LUCIFER:
UT 2009 October 30

Observer: DThompson (@LBT)
Telescope Operator: ACeranski
SW support: CBiddick (on call)

Summary

  • Mirror 6C warmer than outside at beginning of the night
  • All-sky monitor iced over. Doesn't this thing have a heater?
  • UT11:13 opened dome
    • Presets and offsets work as expected
    • Initial collimation strongly comatic
    • Collimation model looks OK 84 > EL > 26
    • Pointing looks off from current model by enough to require frequent re-pointings
    • M1 down to -8.3C
  • 13:16 Closed dome (2 hours)

Preparations

Details

UT 00:40 around sunset. Toutside=-10.4 Tmirror=-4.4. We have been ventilating the dome and mirror since ~23UT. At sunset ~50% cloud cover but it seems to be dissipating quickly.

00:47 Startup of LUCIFER UI. No problems.

00:54 Temps dropping (-10.9C now) and humidity rising (85%). Dome is closed but we are continuing to cool the mirror as long as possible.

00:59 I am keeping the target directory as /data/luci/20091029 to not mess up the LUCIFER team's convention of using local dates (but I recommend switching to UT as soon as possible). Next file correctly set to 0041. Manual read of 1x3.0s o2dcr (0041) saved, looks normal. Because of the rising humidity and warm mirror we are keeping the building facing the wind so the rear vent doors can be opened for cooling. Drives are off, so no preset test yet.

01:05 There appears to be some minor amount of frost on the dome of the all-sky camera. And by 01:44 it is completely covered.

01:46 Execute darks.txt script. Not getting telescope data. Takes (all o2dcr):
  • 2 1x2.5s 0042, 0043
  • 2 1x5.0s 0044, 0045
  • 2 1x25.0s 0046, 0047
  • 2 1x125.0s 0048, 0049
  • 2 1x600s 0050, 0051
  • 2 1x1200s 0052, 0053

Repeat script (merdarks) but with mer (min expo 10s):
  • 2 1x10.0s 0054, 0055
  • 2 1x15.0s 0056, 0057
  • 2 1x25.0s 0058, 0059
  • 2 1x125.0s 0060, 0061
  • 2 1x600s 0062, 0063
  • 2 1x1200s 0064, 0065

04:14 Switch to K+clear (thermal in dome light). 0066+0067 test exposures. Kdome.txt takes o2dcr data:
  • 2 1x2s 0068, 0069
  • 2 1x3s 0070, 0071
  • 2 1x2s 0072, 0073
  • 2 1x5s 0074, 0075
  • 2 1x2s 0076, 0077
  • 2 1x8s 0078, 0079
  • 2 1x2s 0080, 0081
  • 2 1x12s 0082, 0083
  • 2 1x2s 0084, 0085
  • 2 1x20s 0086, 0087
  • 2 1x2s 0088, 0089
  • 2 1x30s 0090, 0091
  • 2 1x2s 0092, 0093
  • 2 1x40s 0094, 0095
  • 2 1x2s 0096, 0097
  • 2 1x50s 0098, 0099
  • 2 1x2s 0100, 0101
  • 2 1x55s 0102, 0103
  • 2 1x2s 0104, 0105
  • 2 1x60s 0106, 0107 (saturated in center)
  • 2 1x2s 0108, 0109
  • repeated o2dcr 0110-0151 (UT 04:36)
  • repeated mer (starting at 10s) merKdomes.txt 10s-60s files: 152-185
  • repeated mer (starting at 10s) merKdomes.txt 10s-60s files: 186-219

05:40 insert blind-blind filters. RH98.3%, -11.3 outside, -4.2C M1. All-sky camera completely frosted over...doesn't this thing have a heater? Clouds blowing over, heavily overcast.

09:44 Ventilating again. Humidity has dropped enough to crack a vent door, but there is ice on the railings outside.

11:13 Believe it or not, we are opening! Toutside=-10.3 Tmirror=-6.1 RH=68%and dropping still.

11:19 WT10_250 to correct pointing. Source close to center

11:36 BS9118_GS96 ACTIVE on-axis on guide star to collimate. We are collimated with seeing varying from 1.5 to 3.0 arcsec!

11:49 BS9118_GS96 TRACK/PARALLACTIC to re-correct pointing

11:55 Recorrect pointing here IE=-73.8 CA=-56.1

Note for Marcus: the "set telescope" button seems to unset itself inappropriately.

11:57 BS9118_GS96 at PA=150, ACTIVE mode...GS is a 3" double but GCS selected the marginally brighter one. It should not be a problem for the WFS. FWHM=1.4" on guider

12:02 execute dither1.txt. Does "jitter 10 90.0" on BS9118. 6x10s integrated o2dcr at each position. Oops...K+blind used, not K+clear. Files 0228 - 0230 (script "exit"ed).

12:08 execute dither1.txt (K+clear)
Probe is failing to move to position...wrong PA? Script "exit"ed. Files 0231-0233. Verified that PA=150 is correct, the GS is ~160 arcsec off axis from BS9118.

12:16 Try it again after homing XY stages on AGw...dither1.txt. Needed to send preset first.

12:19 And again...dither1.txt. dither positions: 0.0,0.0 3.6,-19.1 18.1,-39.8 23.5,-43.9 -6.1,8.5 12.2,14.4 -24.7,36.6 -40.3,-11.4 25.9,-40.5 26.2,17.0 Files 0234-0243

Note: there is strong pickup noise on the acquisition image.

Quick check of collimation model (TRACK mode, 10s exposures at each position). Started collimated on BS9118 at EL~51. C07+29 0244 EL84 C07+11 0245 EL66 C07-10 0246 EL46 C07-30 0247 EL26 C07+06 0248 EL63 ACTIVE. First WFS rms=1425
Collimation seems OK after start-of-night correction

Quick check of pointing (TRACK mode, BrG): WT10_235 0248 WT10_269 0249 WT10_346 0250 WT10_388 0251 WT10_314 0252
Pointing seems OK (better than expected), but only checked a few stars to the west. You will have to correct often, the offsets seen are ~30 arcsec, enough to put the guide star off the acquisition frame.

13:14 BS9126 gs=19 PA=202 ACTIVE after making 32" change in CA, 9 in IE (undoing some of the earlier corrections). Preset failed, strong gradient on acquisition image, too bright for guiding anymore.

13:16 Close dome. Insert blind-blind filters in LUCIFER.

Observing Plan

Closed dome
  • Test writing LUCIFER data to archive1 /newdata disk
  • Verify ingestion into archive, propagation to Tucson and on
  • Familiarization with LUCIFER UI

* Correct initial collimation/pointing with IE/CA

* Run dice5 pattern(s) to check that things still work as in May/June

* Check open-loop tracking performance (rms on off-axis GS vs. rms on LUCIFER) as reference data for comparison to post-Nov-HBS work and this has not been checked in a while.

* Collect astrometric field data to check telescope scale at colder temps

* Check guide star FWHM/rms vs LUCIFER image quality (closed loop DIQ)

* Make new collimation model, implement (optional if it looks good)
  • Derive and implement before pointing model obs

* Make new pointing model, implement (if coll model good)
  • Have TO start the pointing log!
  • Check first position in AZ > 360 degrees wrap
  • Use ARC method to measure AGw/Rotator center
  • 28 points are sufficient, more is always better
  • Can use NB filter in LUCIFER, the pointing stars are ~7.5 mag
  • needs stable temps!
  • May need to do several short PMs as temps are changing rapidly

* Take Field aberration measurements if seeing is good (time-series, stability measurements)

* Demonstrate guiding to 18th mag (Tel=GUIDE)
* WFS on 18th mag. Will likely need to set the wfs max exposure time to 120s
* Quantify the guide/wfs limits with mag and seeing vs. LUCIFER DIQ

-- DavidThompson - 30 Oct 2009
Topic revision: r6 - 30 Oct 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback